OpenStudy (anonymous):

sin α = 15/17 where π/2 < α < π cos β= − 3/5 where π < β < 3π/2 cos θ= 7/25 where −2π < θ < −3π/2 (a) sin(α + β)= ? (b) cos(α + β)= ? I've tried every possible way to get the answer, but I end up with an incorrect solution.

OpenStudy (anonymous):

need to use trig identities for sin(A+B) and cos(A+B)

OpenStudy (anonymous):

\[\sin \left( \alpha+\beta \right)=\sin \alpha \cos \beta+\cos \alpha \sin \beta \]

OpenStudy (anonymous):

\[\cos \left( \alpha+\beta \right)=\cos \alpha \cos \beta-\sin \alpha \sin \beta \]

OpenStudy (anonymous):

\[\frac{ \pi }{ 2}<\alpha <\pi,\cos \alpha~is~negative.\]

OpenStudy (anonymous):

I tried doing that and I still keep getting the wring answer, I don't know what I'm doing wrong.

OpenStudy (anonymous):

wrong*

OpenStudy (anonymous):

\[\pi <\beta<\frac{ 3 \pi }{ 2 },\sin \beta~is~negative\]

OpenStudy (anonymous):

can you show me your work?

OpenStudy (anonymous):

ok, so I've done (15/17) (-3/5) + (-5/17)(-4/5)= -45/85 + 20/85= -25/85

OpenStudy (anonymous):

I don't know if that's the correct answer?

OpenStudy (anonymous):

would you like help with this

OpenStudy (anonymous):

Yes pleaaaaase.

OpenStudy (anonymous):

sin α = 15/17 where π/2 < α < π this is 2nd quadrant cos β= − 3/5 where π < β < 3π/2 this is third quadrant cos θ= 7/25 where −2π < θ < −3π/2 this is equivalent to (−2π+ 2π) < θ < (−3π/2 + 2π) 0 < θ < π / 2 (a) sin(α + β)= ? (b) cos(α + β)= ?

OpenStudy (anonymous):

are we using cos θ in this problem?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

that was extra?

OpenStudy (anonymous):

basically it was part of the question. The problem mainly focuses on the first two equations.

OpenStudy (anonymous):

use a =α , and b = β sin(a) = 15/17 cos(a) = -8 / 17 <--- show this using unit circle cos (b)= -3/5 sin(b) = -4/5 sin(a + b) = sin(a) cos(b) + cos(a) (sin b) = 15/17 *(-3/5) + (-8)/17 *(-4/5) = -13/85

OpenStudy (anonymous):



OpenStudy (anonymous):

Not exactly a unit circle, lets say a circle with radius 17

OpenStudy (anonymous):

Oh. I found the answer for the second half, it's 84/85. Thank you for your help!

OpenStudy (anonymous):

you can use pythagorean theorem to find the unknown side of that triangle